OpenStudy (anonymous):

Let Y be a continuous random variable with probability density function f(y) = k (2 + 3y), 1

OpenStudy (anonymous):

Let Y be a continuous random variable with probability density function f(y) = k (2 + 3y), 1 <y <3 and 0 elsewhere. Find k and E(1/y)

OpenStudy (amistre64):

\[\int_{1}^{3}(k2+3ky)dy=\left.2ky+\frac{3k}{2}y^2\ \right|_{1}^{3}\] \[6k+\frac{27}{2}k-2k-\frac{3}{2}k=1\] \[4k+\frac{24}{2}k=1\] \[k=\frac{2}{32}=\frac{1}{16}\] if i see it right

OpenStudy (amistre64):

i cant tell what E(1/y) is spose to mean tho. Maybe expected value of 1/y?

OpenStudy (anonymous):

yea i figured out the K value and you are right but didn't really found how to solve for E.... but i did (1/y)f(y) integrated from 1 to 3